You purchased tickets for 5 adults and 2 children. The total cost of all the tickets was $97.00. The child's tickets cost $4.00 less than an adult ticket. What is the cost of an Adult ticket and the cost of the child's ticket?
Answer by maciaemi(11)

a = adult tickets
c = child's tickets
Step 1:
5a + 2c = 97
c = a - 4
Step 2:
5a + 2(a - 4) = 97
5a + 2a - 8 = 97
7a = 105
a = 15
Step 3:
c = a - 4
c = 15 -4
c = 11

adult ticket = $15
child's ticket = $11